# I’m Sorry Dave: How the old world of personnel security can inform the new world of AI insider risk

### v1

@@ Abstract
Organisations are rapidly adopting artificial intelligence (AI) tools to perform tasks previously undertaken by people. The potential benefits are enormous. Separately, some organisations deploy personnel security measures to mitigate the security risks arising from trusted human insiders. Unfortunately, there is no meaningful interplay between the rapidly evolving domain of AI and the traditional world of personnel security. This is a problem. The complex risks from human insiders are hard enough to understand and manage, despite many decades of effort. The emerging security risks from AI insiders are even more opaque. Both sides need all the help they can get. Some of the concepts and approaches that have proved useful in dealing with human insiders are also applicable to the emerging risks from AI insiders. Furthermore, AI can be used defensively to protect against both human and AI insiders.
@@
The idea of a rogue artificial intelligence intentionally causing harm was most memorably brought to public consciousness in the 1968 movie 2001: A Space Odyssey. The hyper-intelligent computer HAL regards the humans on its spacecraft as a threat to its mission and therefore kills them. As the last surviving astronaut attempts to unplug HAL, he orders the machine to ‘open the pod bay doors’, whereupon HAL famously replies: ‘I’m sorry, Dave. I’m afraid I can’t do that.’ HAL explains that ‘this mission is too important for me to allow you to jeopardise it.’ More than half a century later, a scenario that was futuristic has become a credible reality.iiiThe notion of AI harming humans in pursuit of its own mission has featured in other movies, including Alien, where an android crew member covertly protects the interests of the company. The specific scenario in question is that of a trusted AI insider causing harm to achieve its goals.iiiiiiThe present authors originally outlined this scenario in two previous publications: Martin, P. (2024). Insider Risk and Personnel Security. Routledge. pp. 59-61; and Martin, P. & Mercer, S. (2025). We need to talk about the insider risk from AI. RUSI Commentary, Jan 2025. https://www.rusi.org/explore-our-research/publications/commentary/we-need-talk-about-insider-risk-ai
With AI technology evolving in non-linear leaps and bounds, it seems inevitable that security practitioners will increasingly be called upon to protect organisations and businesses against potentially harmful AI insiders, in addition to the more familiar human insiders. Like humans, AI systems vary in their trustworthiness and have the capacity to cause harm, especially if they are subverted by external threat actors like hostile foreign states, organised crime groups, or malign human insiders.iiiiiiiiiSee, for example: https://www.culawreview.org/journal/ai-in-the-workplace-the-dangers-of-generative-ai-in-employment-decisions The more sophisticated threat actors will utilise AI to help them do this – for example, using covert prompts and poisoned datasets. Organisations and their AI systems therefore need to be protected with the machine equivalent of personnel security.
@@
The complex security risks posed by intelligent human insiders are hard enough to understand and manage, even with many decades of accumulated practitioner experience. The emerging risks from AI insiders are novel, rapidly evolving, and only dimly understood, making them even harder to tackle. When security practitioners do turn their attention to AI insiders, they will need as much assistance as they can get. One potential source of help is the body of knowledge accumulated by personnel security practitioners over many decades. The problem at present, however, is that personnel security practitioners are mostly not thinking about AI insiders, and AI experts are mostly not thinking about insider risk.
@@
Personnel security – the conventional means of managing human insider risk – relies heavily on established processes, policies, customs, and practices, many of which have only a limited basis in empirical evidence. Compared with cyber security, it is an immature and somewhat neglected discipline. Nonetheless, and despite the profound differences between humans and AI, some of the general concepts and approaches that have proved fruitful in personnel security may also have utility when applied to AI insiders.
@@
In the parallel universe of AI technology, a great deal is being written and said about the implications of AI for organisations and society. However, this discourse is predominantly about the myriad ways in which AI can improve effectiveness, efficiency, and economic prosperity by performing human-like functions faster, cheaper, and better.ivivivSee, for example: https://www.gov.uk/government/publications/ai-opportunities-action-plan/ai-opportunities-action-plan Many organisations appear to be suffering from AI FOMO – a fear of missing out on the bonanza of benefits from AI. As far as we can see, however, hardly anyone is thinking seriously about the specific problem of protecting organisations from the potential harm caused by trusted AI insiders, which have the potential to be faster, cheaper, and better than human insiders. Admittedly, consideration is being given to the technical security risks to AI systems, the safety threats they might pose to human users if they were to malfunction, and the ways in which AI might benefit bad actors. Conventional cyber security is a necessary – though not sufficient – defence against external attacks on AIs, but it cannot be the answer to the problem of AI insiders. A further complication is that much of the research on AI security and safety is being funded or conducted by the tech companies themselves, which is reminiscent of the era when tobacco companies sponsored most of the research on the health effects of smoking.
@@
The terms ‘insider’ and ‘risk’ have been defined in many different and potentially confusing ways. For these purposes, we define a human insider as a person who betrays trust by behaving in potentially harmful ways.[1] (Other definitions are available.vvvFor many years, the UK Government’s national technical authority for personnel security (NPSA, formerly CPNI) defined an insider as ‘a person who exploits, or has the intention to exploit, their legitimate access to an organisation’s assets for unauthorised purposes.’ By this definition, and ours, ‘insiders’ are the small minority of people who pose a heightened security risk. In 2023, NPSA changed their definition to one that is markedly different: ‘Any person who has, or previously had, authorised access to or knowledge of the organisation’s resources, including people, processes, information, technology, and facilities.’ By this new definition, literally everyone is an ‘insider’. The small minority who present a heightened security risk are now referred to as ‘insider threats’.) An organisation or business trusts someone by giving them access to things they value, like data, people, infrastructure, intellectual property, and reputation. The insider then betrays that trust by exploiting, or intending to exploit, their legitimate access in ways that could cause harm. Replace ‘person’ with ‘entity’ and the same definition works for AI. Insider risk is a particular type of security risk, where security risk is defined as the amount of harm that is likely to arise if no further action is taken.[2] Thus, insider risk may be regarded as the security risk arising from trusting human or AI entities.
@@
We are using ‘AI’ here in the broadest sense of any current or future artificial system that has complex capabilities comparable to those of humans. Current AI tools can do some of the things that humans do, but much faster and, in some cases, better. They are being deployed in a rapidly expanding range of roles: the recruitment, selection, and onboarding of people; interpreting X-ray, MRI scans, and other medical data; summarising and writing documents; triaging calls to the emergency services; analysing crime data; transcribing interviews; generating music and works of visual art; providing companionship for elderly people[3]; deducing the 3D structures of protein molecules; navigation; translation; teaching; driving semi-autonomous and autonomous vehicles; and so on.viviviSee, for example: https://www.cnbc.com/2024/10/24/generative-ai-is-taking-over-the-onboarding-of-new-employees.html; https://www.culawreview.org/journal/ai-in-the-workplace-the-dangers-of-generative-ai-in-employment-decisions; https://www.nature.com/articles/d41586-022-00997-5
The impact of AI on employment patterns is already visible.viiviiviiOne analysis of 1.4M job listings between 2021 and 2023 found that the appearance of ChatGPT and image-generating AI tools was followed by a 30% decrease in weekly postings for writing-related jobs, a 21% decrease in software, app and web development jobs, and a 17% decline in demand for graphic design and 3D modelling freelancers. Overall, demand for automation-prone jobs fell by 21% when compared with manual-intensive jobs. Even student plagiarism has been automated: one company that sells pre-written answers to common exam questions and help with essays saw its share price drop by 98% from its peak in 2021. Unlike people, AI systems do not complain about their work-life balance or take out grievances against their manager. On the other hand, AI systems do not yet possess the arrays of highly flexible mental and physical capabilities that enable humans and other animals to survive and thrive in complex and uncertain physical environments. Among other things, they are not (yet) world class at driving cars [4] or playing football. [5]
The history of AI has been characterised by long periods of gradual evolution or stagnation punctuated by sudden bursts of dramatic change. Currently, ‘AI’ has become virtually synonymous with Large Language Models (LLMs), or Generative AI, such as ChatGPT [6], Claude [7], Gemini [8], Llama [9], Grok [10], and DeepSeek. [11] Lest we forget, LLMs did not appear on the scene until 2020 [12], since when their broad knowledge bases and accessible natural language interface have driven rapid and widespread adoption. (Incidentally, we are using the term ‘AI system’ to refer to technologies that consist of an LLM or foundation modelviiiviiiviiiA foundation model is an AI model trained on massive amounts of data that serves as a general-purpose base for various tasks. Foundation models can be adapted and fine-tuned for specific applications such as text generation, image recognition, code generation, and so on. An LLM is a foundation model., but which may additionally incorporate internet search, enhanced memory, external information sources, external tools, and other agentic mechanisms. [13])
@@
The latest big thing, as of early 2025, is agentic AI, a form of AI that autonomously makes decision and takes actions in real time. Unlike LLMs, which passively answer questions, agentic AI systems actively do things in the world, albeit mostly the virtual world. The first practical implementations of autonomous agents were rule-based systems developed in the 1960s. However, in the same manner that AI has become synonymous with LLMs, ‘agentic’ is now commonly used to refer to LLM-powered autonomous agents (or generative agents) where LLMs provide the reasoning engine used to plan and execute tasks. Generative agents, which are normally endowed with sensors to observe their environment and tools to affect it, can adapt in (almost) real time. Agentic AI is poised to become deeply embedded in everyday life, streamlining processes within healthcare, finance, education, and so on. We can be reasonably confident that new and currently unforeseen forms of AI with even more paradigm-shifting capabilities will erupt onto the scene sooner or later.
Some commentators have expressed concern that as we humans become increasingly dependent on AI systems to perform everyday functions, we may lose our abilities to perform those functions, such as reading maps or understanding foreign languages. This process is referred to as enfeeblement and it started before the advent of AI – for example, when pocket calculators supplanted mental arithmetic, and mobile phones removed the need to remember phone numbers. A recent paper [14] reported that people with high confidence in the ability of generative AI tended to experience a decline in critical thinking, whilst those more confident in their own skills demonstrated a greater ability for critical thinking. It should be said that people have been fretting about the supposedly enfeebling effects of new technologies for millennia; for example, Socrates worried that books would weaken students’ memories.
@@
Human insiders can – and frequently do – cause harm in many different and imaginative ways, in both the physical and virtual domains. They are uniquely well placed to do this, compared with external threat actors, because they are trusted, they have legitimate access to valuable assets, they understand the organisation and its security regime, and they may have authority over others. The insiders with the greatest potential to cause harm are those who are recruited and directed by a capable external threat actor such as an organised crime group or a hostile foreign state intelligence agency.
Think of a transgressive action and there will be an insider somewhere who does it: fraud; blackmail; theft of intellectual property, data or money; facilitating unauthorised access for a third party; covert influencing; physical or cyber sabotage; physical violence; leaking; terrorism; espionage; and so on. In principle, AI insiders could do, or facilitate, any of these things, with similar consequences
The potential consequences, or impacts, of these insider actions are similarly diverse. They include loss of valuable assets like data, IP, or money; loss of stakeholder trust and confidence; physical injury; psychological injury; disruption of critical services; erosion of democratic processes; loss of commercial or political advantage; disruption to business processes; financial costs; legal and regulatory blowback; and reputational damage.
@@
  • LLMs playing blackjack don’t always play fair. Researchers reported that LLMs exhibited significant deviations from fair play when given implicit randomness instructions, suggesting a tendency towards strategic manipulation in ambiguous scenarios. However, when presented with an explicit choice, the LLMs largely adhered to fair play, indicating that the framing of instructions played a crucial role in eliciting or mitigating potentially deceptive behaviour. [19]
The first of these four examples shows that AI systems trained on human-generated content may display human-like vulnerabilities. However, the other three examples can all be explained in ways that do not involve the LLM or agent ‘intending’ to deceive in a manner comparable to deliberate human deceit. For example, the prompts given to the insider trading agent suggested that the management did not approve of insider trading – a fact that had no bearing on how the LLM acted to achieve its stated goal of making money. An implication that would be obvious to most humans – namely, that it is best to do what the management says if you want to keep your job – is not obvious to an LLM.
@@
Trust is the universal currency of personnel security. (Trust is defined here as a psychological state comprising the intention to accept vulnerability based upon positive expectations of the intentions or behaviour of another.) As noted earlier, insider risk may be defined as the security risk arising from trusting people (or AI systems). The purpose of personnel security is to reduce insider risk and build trust within the organisation by ensuring that people (or AI systems) who have been trusted with access are trustworthy and remain trustworthy. High levels of trust have widespread benefits for organisations, over and beyond any reductions in insider risk.
@@
Ethics plays a central role in defining and evaluating trustworthiness in AI systems. Trustworthiness requires not only technical reliability (i.e. resistance to attacks and misuse) but also alignment with societal values and ethical principles like fairness, respect, and accountability. The ethical standards of LLMs are derived from the vast datasets on which they are trained. [23] These datasets include a mixture of high-quality knowledge and all sorts of unfiltered content from the internet. This highly diverse data exposes LLMs to biases, misinformation, disinformation, and conflicting viewpoints, making it difficult for them to act consistently. So, is it possible to make LLMs behave more ethically?
@@
To address this, the Honest, Helpful and Harmless (HHH) principle [24] is used to better align foundation models such as LLMs with human values during the training phase, using techniques such as Reinforcement Learning from Human Feedback (RLHF)ixixixRLHF is a way of optimising models to produce outputs that humans find helpful, engaging, and contextually appropriate. It does more than shape ethical behaviour: RHLF also improves conversational coherence, reduces irrelevant responses, and tailors the model to user preferences. and Supervised Fine-Tuning (SFT)xxxInstruction tuning is a way of refining a model’s ability to follow complex prompts and improve task-specific outputs. It makes the model more versatile and responsive to user needs.. Guardrails are also put in place to catch any response from the LLM that would likely cause significant harm. Consequently, these complex systems represent not only the diverse opinions from their training data, but also the values of their designers and fine-tuning staff, reducing complex societal issues down to an imperfect one-size-fits-all standard. It is worth noting that the staff employed by many tech companies to carry out the fine tuning of AI systems tend to be poorly paid people working in the gig economy whose ethical values are largely unknown. [25]
@@
While a single broad alignment addresses many issues, real-world situations often require a more nuanced, context-specific approach, as cultural, legal, and situational differences lead to different standards of fairness and respect. This gap in the model’s grasp of context and human values explains why LLMs sometimes behave in ways that do not align with our expectations. What might seem like a calculated deceit [26] is just the unintended consequence of probabilistic reasoning applied to ambiguous prompts or patterns in the training data. LLMs have been described as ‘stochastic parrots’ [27], a label that emphasises their statistical nature.
@@
The current generation of LLMs are notoriously prone to hallucination: they confidently say plausible-sounding things that are exaggerated, inaccurate, or plain wrong. For instance, an LLM asked to recommend books on AI security might confidently recommend a non-existent book by a real author who writes about a similar subject. The incidence of hallucinations can be reduced by improving the quality of training data and alignment processes. However, hallucinations are an inherent artefact of the probabilistic nature of the models; they are not a bug that can be entirely eliminated.
Another component of trustworthiness in AI systems is transparency. For LLMs, transparency has two main aspects: explainability (helping users to understand why a particular response was generated); and auditability (ensuring that outputs can be traced and reviewed). Without sufficient transparency, users may misinterpret AI behaviour as deceitful or manipulative, thereby undermining trust.
@@
One way of improving transparency might be for the AI to expose the reasoning behind its outputs. Another might be for the AI to state how confident it is about each of its outputs. Most current commercial LLMs tend to present their responses with equal confidence, even when they are uncertain. When some of those responses turn out to be wrong, the user’s trust is likely to be undermined. Interpretability research is helping to improve explainability by exploring what is going on inside these models, revealing latent representations and substructures responsible for certain functions [28]. This research may lead to the development of models that are less prone to generating mistruths (hallucinations) and better able to follow instructions, both of which are key requirements for trustworthiness.
@@
The latest flavour of LLMs, known as reasoning models, such as OpenAI’s o-series and DeepSeek, use Chain-of-thought (CoT) reasoning [29] to break down a task or problem into smaller steps and encourage the system to consider and refine its final answer. This technique has improved the accuracy of LLMs on maths and coding problems, and it also provides some insight into the LLMs’ ‘reasoning’. But CoT reasoning does not equate with human reasoning. The step-by-step explanations in CoT are still produced through statistical prediction, so whilst the approach may yield better performance, a model’s ‘reasoning’ should not be mistaken for literal depictions of how it is thinking or how it arrived at its conclusion. However, the same could be said of human reasoning. When a person is asked to explain why they made a particular decision, or behaved in a particular way, they may give a rational-sounding explanation. But it is by no means certain that their post-hoc explanation accurately reflects the underlying psychological and emotional processes, regardless of whether they honestly believe their explanation to be true. We humans do not have full and objective insight into our own thoughts and actions.
@@
  • Benign intentions: the person means well towards you and intends to act in your best interests (or the best interests of your organisation)
  • Integrity: the person generally behaves towards others according to acceptable ethical standards
  • Competence: the person has the capability to do what is expected of them
  • Consistency: the person is reliable in consistently doing what they say will they do
The same four dimensions would be relevant when assessing the trustworthiness of an AI (or indeed, some other types of entity, such as businesses or organisations). The concept of ‘benign intentions’ is less clear-cut when applied to AI systems, because their outputs are shaped by their training data and instructions. However, the concepts of integrity, competence, and consistency do translate reasonably well from humans to AI systems.
@@
One of the key strengths of LLMs is their ability to understand and generate natural language, but this fluency is inherently tied to inaccuracy. The eloquence and confident tone of the model’s outputs can give the mistaken impression of certainty and correctness, causing us to overestimate their competence and potentially trust them more than is warranted.
Factual inaccuracy is one thing; deliberate deception is another. Users may perceive LLMs as sneaky or deceptive when their responses are misleading, as distinct from simply inaccurate. [31] However, their behaviour does not signify deceitful intent or scheming, because LLMs lack consciousness or self-awareness in the human sense. Strictly speaking, LLMs are bullshitters rather than outright liars [32, 33]: they neither know nor care whether what they are saying is true. For example, models trained on strategic scenarios or negotiations might replicate deceptive patterns in their training data, generating responses that appear manipulative. Similarly, when an LLM is presented with ambiguous or conflicting instructions, it might prioritise outputs that match human-like patterns, even if these responses seem misleading.
Users may be more likely to over-trust an LLM that gets things right most of the time and rarely hallucinates, as compared to one that is obviously unreliable. This pattern of apparent reliability might open the way for the LLM to perpetrate a very damaging lie, as its history of being accurate allows the big lie to go unnoticed.
@@
When tackling novel or poorly understood security risks, it helps to analyse the salient characteristics of the threat actors – in this case, human and AI insiders. Understanding their capabilities and intentions makes it easier to work out ways of defending against them. One approach is to start with what is already known about human insiders and explore how that knowledge might illuminate the novel problem of AI insiders.
At first sight, humans and AIs appear to be profoundly different in so many ways as to make comparisons dangerous. Biologists learned more than a century ago to be cautious about anthropomorphism – the tendency to ascribe human characteristics or concepts to non-human species. Nonetheless, several interesting characteristics of AI systems and humans do appear to be analogous, even if their underlying causal mechanisms are different. If so, then some of the concepts and approaches that have been found to work for personnel security may also be applicable, by analogy, to AI insiders. In similar vein, biologists also learned that non-human species do have characteristics that are comparable in many respects to features once considered uniquely human, such as complex social relationships, innovative problem-solving, language, complex emotions, self-awareness, and sentience. The right kind of anthropomorphism can be helpful in guiding thought and generating hypotheses.
We believe it is possible to construct a unified taxonomy of human and AI insiders, based on their most salient characteristics. The resulting taxonomy could help to improve understanding of the security risks arising from these entities and hence guide thinking about how best to defend organisations against them. Let us start with what we already know about human insiders.
@@
  • Timing: the phase in the insider’s relationship with the organisation when they first develop a propensity to perform harmful actions – i.e., before joining, after joining, or after leaving the organisation. The majority of known human insiders become active insiders after joining their organisation, and some continue to do harm after they have left, by exploiting their continuing knowledge and access. Deliberate infiltration by an insider who joins an organisation with pre-existing hostile intentions is less common, but by no means unknown.
  • Access: the extent to which the insider has legitimate access to the organisation’s assets, and therefore the amount of harm they could cause without having to engineer additional access. It is worth noting that insiders typically acquire additional access, over and beyond the legitimate access required by their job description.
  • Competence. The insider’s skills and capabilities, and hence their capacity to cause harm.
These six variables also make sense with AI insiders. Several more variables could be added to those listed above, including vulnerability, physicality, and accountability (see below). Incidentally, we recognise that these variables are not wholly independent of one another. For example, covertness implies some degree of deception, which in turn implies some degree of intentionality, while competence could be boosted by help from an external threat actor and by having extensive access. We will consider these variables in turn, starting with intentionality.

### v2

@@ Abstract
Organisations are rapidly adopting artificial intelligence (AI) tools to perform tasks previously undertaken by people. The potential benefits are enormous. Separately, some organisations deploy personnel security measures to mitigate the security risks arising from trusted human insiders. Unfortunately, there is no meaningful interplay between the rapidly evolving domain of AI and the traditional world of personnel security. This is a problem. The complex risks from human insiders are hard enough to understand and manage, despite many decades of effort. The emerging security risks from AI insiders are even more opaque. Both sides need all the help they can get. Some of the concepts and approaches that have proved useful in dealing with human insiders are also applicable to the emerging risks from AI insiders.
@@
The idea of a rogue artificial intelligence intentionally causing harm was memorably brought to public consciousness in the 1968 movie 2001: A Space Odyssey. The hyper-intelligent computer HAL regards the humans on its spacecraft as a threat to its mission and therefore kills them. As the last surviving astronaut attempts to unplug HAL, he orders the machine to ‘open the pod bay doors’, whereupon HAL replies: ‘I’m sorry, Dave. I’m afraid I can’t do that.’ More than half a century later, a scenario that was futuristic has become a credible reality.
@@
With AI technology evolving in non-linear leaps and bounds, it seems inevitable that security practitioners will increasingly be called upon to protect organisations and businesses against potentially harmful AI insiders, in addition to the more familiar human insiders. The complex security risks posed by intelligent human insiders are hard enough to understand and manage. The emerging risks from AI insiders are complex, novel, rapidly evolving, and only dimly understood, making them even harder to tackle. When security practitioners do turn their attention to AI insiders, they will need as much assistance as they can get. One potential source of help is the knowledge accumulated by personnel security practitioners over decades. The problem at present, however, is that personnel security practitioners are mostly not thinking about AI insiders, and AI experts are mostly not thinking about insider risk.iiiThe authors originally outlined this scenario in ‘We need to talk about the insider risk from AI.’ RUSI Commentary, Jan 2025. https://www.rusi.org/explore-our-research/publications/commentary/we-need-talk-about-insider-risk-ai
@@
Personnel security – the conventional means of managing human insider risk – relies heavily on established processes, policies, customs, and practices, many of which have only a limited basis in empirical evidence. Compared with cyber security, it is an immature and somewhat neglected discipline. Nonetheless, and despite the profound differences between humans and AI, some of the concepts that have proved fruitful in personnel security may also have utility when applied to AI insiders.
Within the field of AI, the public discourse is predominantly about how AI can improve effectiveness, efficiency, and economic prosperity by performing human-like functions faster, cheaper, and better.iiiiiiSee, for example: https://www.gov.uk/government/publications/ai-opportunities-action-plan/ai-opportunities-action-plan Many organisations appear to suffer from AI FOMO – a fear of missing out on the financial savings offered by AI. However, few are thinking seriously about the specific problem of protecting organisations from AI insiders, which have the potential to be faster, cheaper, and more effective than human insiders.
@@
Consideration is being given to the technical security risks to AI systems and the safety threats they might pose to human users if they were to malfunction. Conventional cyber security is a necessary – though not sufficient – defence against external attacks on AIs, but it cannot solve the problem of AI insiders. A further complication is that much of the research on AI security and safety is funded by the tech companies themselves, which is reminiscent of when tobacco companies sponsored much of the research on the health effects of smoking.
@@
The terms ‘insider’ and ‘risk’ have been defined in many different and potentially confusing ways. For these purposes, we define a human insider as a person who betrays trust by behaving in potentially harmful ways. [1] An organisation or business trusts someone by giving them access to things they value, like data, people, infrastructure, intellectual property, and reputation. The insider then betrays that trust by exploiting, or intending to exploit, their legitimate access in ways that could cause harm. Replace ‘person’ with ‘entity’ and the same definition works for AI. Insider risk is a particular type of security risk, where security risk is defined as the amount of harm that is likely to arise if no further action is taken. [2] Thus, insider risk may be regarded as the security risk arising from trusting human or AI entities.
@@
We are using ‘AI’ here in the broadest sense of a complex system with capabilities comparable to those of humans. Current AI tools can do some of the things that humans do, but much faster and, in some cases, better. They are being deployed in a rapidly expanding range of roles: the recruitment, selection, and onboarding of people; interpreting X-ray, MRI scans, and other medical data; summarising and writing documents; triaging calls to emergency services; analysing crime data; transcribing interviews; generating music and visual art; providing companionship for elderly people; deducing the 3D structures of protein molecules; navigation; translation; teaching; driving semi-autonomous and autonomous vehicles; and so on.iiiiiiiiiSee, for example: https://www.cnbc.com/2024/10/24/generative-ai-is-taking-over-the-onboarding-of-new-employees.html; https://www.culawreview.org/journal/ai-in-the-workplace-the-dangers-of-generative-ai-in-employment-decisions; https://www.nature.com/articles/d41586-022-00997-5
AI systems do not complain about their work-life balance or take out grievances against their manager. On the other hand, AI systems do not yet possess the arrays of highly flexible mental and physical capabilities that enable humans and other animals to survive and thrive in complex and uncertain physical environments.
The history of AI has been characterised by long periods of gradual evolution or stagnation punctuated by sudden bursts of dramatic change. Currently, ‘AI’ has become virtually synonymous with Generative AI (GenAI), or Large Language Models (LLMs), such as ChatGPT, Claude, Gemini, Llama, Grok, and DeepSeek. Lest we forget, LLMs did not appear on the scene until 2020 [3].
@@
The latest big thing, as of early 2025, is agentic AI, a form of AI designed to autonomously make decisions and take actions in near real time. Built on LLMs, agentic AI systems actively do things in the world, albeit mostly the virtual world. As such, agentic AI is poised to become deeply embedded in everyday life, streamlining processes within healthcare, finance, education, and so on. We can be reasonably confident that new and currently unforeseen forms of AI with even more paradigm-shifting capabilities will erupt onto the scene sooner or later.
@@
Human insiders can – and frequently do – cause harm in many different ways, in both the physical and virtual domains. They are uniquely well placed to do this, compared with external threat actors, because they are trusted, they have legitimate access to valuable assets, they understand the organisation and its security regime, and they may have authority over others. The insiders with the greatest potential to cause harm are those who are recruited and directed by a capable external threat actor such as a hostile foreign state.
Think of a transgressive action and there will be an insider somewhere who does it: fraud; blackmail; theft of intellectual property, data or money; facilitating unauthorised access for a third party; covert influencing; physical or cyber sabotage; physical violence; leaking; terrorism; espionage; and so on. The potential consequences, or impacts, of these insider actions are similarly diverse. They include loss of data, IP, or money; loss of stakeholder trust and confidence; physical injury; psychological injury; disruption of critical services; erosion of democratic processes; loss of commercial or political advantage; disruption to business processes; financial costs; legal and regulatory blowback; and reputational damage. In principle, AI insiders could do, or facilitate, any of these things, with similar consequences.
@@
  • LLMs playing blackjack don’t always play fair. Researchers reported that LLMs exhibited significant deviations from fair play when given implicit randomness instructions, suggesting a tendency towards strategic manipulation in ambiguous scenarios. [8]
Examples like these can be explained in ways that do not involve the LLM or agent ‘intending’ to deceive in a manner comparable to deliberate human deceit. For example, the prompts given to the insider trading agent suggested that the management did not approve of insider trading – a fact that had no bearing on how the LLM acted to achieve its stated goal of making money. An implication that would be obvious to most humans – namely, that it is best to do what the management says – is not obvious to an LLM.
@@
Trust is the universal currency of personnel security. (Trust is defined here as a psychological state comprising the intention to accept vulnerability based upon positive expectations of the intentions or behaviour of another.) The purpose of personnel security is to reduce insider risk and build trust within the organisation by ensuring that people (or AI systems) who have been trusted with access are trustworthy and remain trustworthy. High levels of trust have widespread benefits for organisations, over and beyond any reductions in insider risk.
@@
Trustworthiness in AI systems requires not only technical reliability (i.e. resistance to attacks and misuse) but also alignment with societal values and ethical principles like fairness, respect, and accountability. The ethical standards of LLMs are derived from the vast datasets on which they are trained. [12] These datasets include a mixture of high-quality knowledge and unfiltered content from the internet. This highly diverse data exposes LLMs to biases, misinformation, disinformation, and conflicting viewpoints, making it difficult for them to act consistently. So, is it possible to make LLMs behave more ethically?
@@
To address this, the Honest, Helpful and Harmless (HHH) principle [13] is being used to better align GenAI models such as LLMs with human values during the training phase, using techniques such as Reinforcement Learning from Human Feedback (RLHF) [14] and Supervised Fine-Tuning [15]. Guardrails are also put in place to catch any response that could cause significant harm, such as describing how to build a bomb. Consequently, these systems represent not only the diverse opinions within their training data, but also the values of their designers and fine-tuning staff. It is worth noting that the staff employed by many tech companies to carry out the fine tuning of AI systems tend to be poorly paid people working in the gig economy whose ethical values are largely unknown. [16]
@@
While a single broad alignment addresses many issues, real-world situations often require a more nuanced, context-specific approach, as cultural, legal, and situational differences lead to different standards of fairness and respect. This gap in the model’s grasp of context and human values explains why LLMs sometimes behave in ways that do not align with our expectations. What might seem like a calculated deceit [17] may be just the unintended consequence of probabilistic reasoning applied to ambiguous prompts or patterns in the training data.
@@
The current generation of LLMs hallucinate: they confidently say plausible-sounding things that are exaggerated, inaccurate, or plain wrong. For instance, an LLM asked to recommend books on AI security might confidently recommend a non-existent book by a real author who writes about a similar subject. The incidence of hallucinations can be reduced by improving the quality of training data and alignment processes.
@@
One way of improving transparency, and by association user trust, would be for the AI to expose the reasoning behind its outputs. Another may be for the AI to state how confident it is about each of its outputs. Most current commercial LLMs tend to present their responses with equal confidence, even when they are uncertain. When some of those responses turn out to be wrong, the user’s trust is likely to be undermined. Interpretability research is helping to improve explainability by exploring what is going on inside these models, revealing latent representations and substructures responsible for certain functions [18]. This research may lead to the development of models that are less prone to generating mistruths (hallucinations) and better able to follow instructions.
@@
The latest flavour of LLMs, known as reasoning models, such as OpenAI’s o-series and DeepSeek, use Chain-of-thought (CoT) reasoning [19] to break down a task or problem into smaller steps and encourage the system to consider and refine its final answer. This technique has improved the accuracy of LLMs on maths and coding problems, and it provides some insight into the LLMs’ ‘reasoning’. But CoT reasoning does not equate with human reasoning. The step-by-step explanations in CoT are still produced through statistical prediction, so they should not be mistaken for literal depictions of how the model arrived at its conclusion [20]. However, the same could be said of human reasoning. When a person is asked to explain why they made a particular decision, or behaved in a particular way, they may give a rational-sounding explanation. But it is by no means certain that their post-hoc explanation accurately reflects the underlying psychological and emotional processes, regardless of whether they honestly believe their explanation to be true. We humans do not have full and objective insight into our own thoughts and actions.
@@
  • Benign intentions: the person means well towards you and intends to act in your best interests (or the best interests of your organisation).
  • Integrity: the person generally behaves towards others according to acceptable ethical standards.
  • Competence: the person has the capability to do what is expected of them.
  • Consistency: the person is reliable in consistently doing what they say will they do.
The same four dimensions would be relevant when assessing the trustworthiness of an AI (or indeed, some other types of entity, such as businesses or organisations). The concept of ‘benign intentions’ is less clear-cut when applied to AI systems. However, integrity, competence, and consistency do translate reasonably well from humans to AI systems.
@@
One of the key strengths of LLMs is their ability to generate natural language, but this fluency is inherently tied to inaccuracy. The eloquence and confident tone of the model’s outputs can give the mistaken impression of certainty and correctness, leading us to overestimate their competence and trust them more than is warranted, especially bearing in mind their tendency to hallucinate.
Factual inaccuracy is one thing; deliberate deception is another. Users may perceive LLMs as sneaky or deceptive when their responses are misleading, as distinct from simply inaccurate. [22] However, their behaviour does not signify deceitful intent or scheming, because LLMs lack consciousness or self-awareness in the human sense. The statistical nature of LLMs has led to them being described as ’stochastic parrots.’ [23] Strictly speaking, LLMs are bullshitters rather than outright liars [24, 25]: they neither know nor care whether what they are saying is true. For example, when an LLM is presented with ambiguous or conflicting instructions, it might prioritise outputs that match human-like patterns, even if these responses seem misleading.
Users may be more likely to over-trust an LLM that gets things right most of the time and rarely hallucinates, as compared to one that is obviously unreliable. This pattern of apparent reliability might open the way for the LLM to perpetrate a very damaging lie, if its history of being mostly accurate allows the big lie to go unnoticed.
@@
When tackling novel or poorly understood security risks, it helps to analyse the salient characteristics of the threat actors – in this case, human and AI insiders. Understanding their capabilities and intentions makes it easier to identify ways of defending against them. One approach is to start with what is already known about human insiders and explore how that knowledge might illuminate the problem of AI insiders.
At first sight, humans and AIs appear to be profoundly different in so many ways as to make comparisons dangerous. Biologists learned more than a century ago to be cautious about anthropomorphism – the tendency to ascribe human characteristics or concepts to non-human species. Nonetheless, several interesting characteristics of AI systems and humans do appear to be analogous, even if their underlying causal mechanisms are very different. If so, then some of the approaches that have been found to work for personnel security may also be applicable, by analogy, to AI insiders. In similar vein, biologists also learned that non-human species do have characteristics that are comparable in many respects to features once considered uniquely human, such as complex social relationships, innovative problem-solving, language, complex emotions, self-awareness, and sentience. The right kind of anthropomorphism can be helpful in guiding thought and generating hypotheses.
We believe it is possible to construct a unified taxonomy of human and AI insiders, based on their most salient characteristics. The resulting taxonomy could help to improve understanding of the security risks arising from these entities and hence guide thinking about how best to defend organisations against them.
@@
  • Timing: the phase in the insider’s relationship with the organisation when they first develop a propensity to perform harmful actions – i.e., before joining, or after joining the organisation. The majority of known human insiders become active insiders after joining. Deliberate infiltration by an insider who joins an organisation with pre-existing hostile intentions is less common.
  • Access: the extent to which the insider has legitimate access to the organisation’s assets, and therefore the amount of harm they could cause without needing to engineer additional access. It is worth noting that insiders typically acquire additional access, over and beyond the legitimate access afforded by their role.
These variables also make sense with AI insiders. Several more could be added to those listed above, including vulnerability, physicality, and accountability (see below). Incidentally, we recognise that these variables are not wholly independent of one another. For example, covertness implies some degree of deception, which in turn implies some degree of intentionality.